### v2

# Elongated vortex quantum droplets in binary Bose-Einstein condensates

Abstract
Stability of elongated (“slender”) quantum droplets (QDs) with embedded unitary and multiple vorticity is a problem that was not solved previously. In this work, we propose a solution which relies upon the use of the spatial modulation of the inter-species scattering length in the binary Bose-Einstein condensates, in the form of a two-dimensional axisymmetric Gaussian, shaped by means of the optical Feshbach resonance. The corresponding effective nonlinear trapping potential supports completely stable elongated QDs with vorticity $S=0$ and partly stable families of elongated QDs with $S=1,2,3,4$ (other nonlinear systems do not maintain stability of vortex droplets with $\geq 2$). We systematically analyze effects of the amplitude and width of the Gaussian modulation, as well as the total number of atoms, on the shape and stability of the QDs, some effects being explained analytically. Collisions between identical QDs with $S=1$ moving in opposite directions along the central axis leads to their merger into still more elongated breathing QDs with the same vorticity, while collisions between QDs with $S=\pm 1$ are quasi-elastic. Moving modulation profiles are able to adiabatically rotate the trapped elongated QDs. Application of a torque to the vector QD sets in the gyroscopic regime of robust precession, which realizes a macroscopic spin-orbit-coupling effect.
I Introduction
Vortices with an elongated three-dimensional (3D) structure have drawn much interest in astrophysics McWilliams1994a ; astro and geophysics McWilliams1994b , plasmas plasma1 ; plasma2 ; plasma3 , superconductors Dolan1989 ; Odobesko2020 ; Hung2012 ; Olszewski2018 , optics Freund1997 ; Sharma2011 , and Bose-Einstein condensates (BECs) Stockhofe2011 ; Liu2013 ; McEndoo2010 ; OKtel2004 ; Engels2002 ; Tikhonenkov2008 ; Tan2023 . However, the realization of elongated self-localized vortices - those stabilized solely by intrinsic interactions - has remained elusive. Quantum droplets (QDs) LHY ; Petrov2015 ; Petrov2016 ; Luca1 ; Jrgensen2018 ; Cabrera2018 ; Ferrier-Barbut2018 ; Cheiney2018 ; Inguscio2018 ; Luca2 ; Yin2021 ; dlw2021 ; Luo2021 ; Cui2025 offer a promising platform to address this challenge, as they inherently support the formation of self-localized vortices without external linear confinement, stable 2D and 3D QDs with embedded vorticity and more complex vortex patterns have been predicted Kartashov2018 ; QDvortex2D ; Tengstrand2019 ; Chen2024 ; Li2024FOPview .
The most intuitive solution lies in dipolar QDs where the stabilization emerges from the competitive balance between anisotropic dipole-dipole interactions (DDIs) Stuhler2005 ; Lahaye2009 and beyond mean-field effects. Although such systems exhibit intrinsic anisotropy Ferrier-Barbut2016 ; Schmitt2016 ; Santos ; Wang2020 ; Bottcher2021 ; Yin2022 , elongated vortex structures have proven fundamentally unstable Cidrim2018 , permitting only short-axis-aligned vortex configurations which the vorticity axis is perpendicular to the direction of the dipolar polarization Li2024PRL ; Li2024FOP .
An alternative approach within quantum droplets emerges from two-component atomic systems, which can sustain stable vortices but their inherent isotropic interactions prevent spontaneous elongation Kartashov2018 . This isotropic interaction is partly due to the dual-component stabilization mechanism, which utilizes magnetic Feshbach resonance (MFR) with dc magnetic field which alters the s-wave scattering length $a_{s}$ of the inter-atomic interaction to achieve uniform modulation of the interaction Tiesinga1993 ; Moerdijk1995 ; Courteille1998 ; Bertozzi1998 ; Inouye1998 ; Timmermans1999 ; Duine2004 ; kohler2006 ; Jones2006 ; Randy ; Chin2010 . In stark contrast, the characteristic feature of optical Feshbach resonance (OFR) - its ability to spatially inhomogeneously modulate scattering lengths Fisher1989 ; Fedichev1996 ; Caputo ; Saito2003 ; Qi2011 ; Chien2012 - reveals an extraordinary possibility for engineering elongated vortex states. It is worth noting that recent advancements in optical manipulation demonstrate OFR’s capacity for micro scale spatial density modulation in quantum droplets Schafer ; Fu , enabling precise morphology design, tailored macroscopic mechanical properties, and novel approaches for droplet dynamics manipulation. This technological progress establishes OFR as a transformative tool for quantum droplet engineering, particularly in realizing elongated vortex configurations.
The objective of the present work is to predict stable vortex QDs with a “slender” (elongated) shape in the binary condensate with the contact interactions, using an effective nonlinear potential RMP (alias pseudopotential pseudo ), that may be induced by a spatially modulated interaction strength. In turn, the modulation can be imposed by means of the spatially inhomogeneous Feshbach resonance - OFR, it can be readily induced by spatially non-uniform optical fields with inhomogeneous intensity $I\left(x,y,z\right)$, which can be used to shape effective nonlinear potentials through the local OFR relation,
$a_{\mathrm{opt}}\varpropto-I\left(x,y,z\right),$ (1)
where $a_{\mathrm{opt}}$ is the OFR-induced change of the scattering length, $\left(x,y,z\right)$ are spatial coordinates, and sign minus implies that the resonant laser illumination induces the inter-atomic attraction force, which can be achieved by adjusting the detuning of the laser Theis2004 ; Bauer2009 ; Yamazaki2010 ; RMP ; Yan2013 .
II Model
In this work, we consider the possibility to use OFR for mapping the laser intensity distribution in the usual Gaussian laser beam, propagating along the $z$ axis, into an effective nonlinear trapping potential with the transverse Gaussian profile, determined as per Eq. (1), viz.,
$a_{\mathrm{opt}}(x,y)=-a^{(0)}\exp\left(-\rho^{2}/L^{2}\right),\ \rho\equiv\sqrt{x^{2}+y^{2}},$ (2)
where $L$ is the beam waist, and $-a^{(0)}<0$ is the amplitude of the OFR-induced modulation of the local scattering length. The objective is to use the axisymmetric profile (2) to stabilize vortex solitons with the angular momentum directed along the $z$ axis and a “slender” (elongated) shape. The setting will be also used to study more sophisticated dynamical regimes, such as gyroscopic motion (precession) of the vortex soliton.
Our study, referring to the use of 39K atoms, implies large spectral detuning between the OFR wavelength and that corresponding to the atomic dipole transition, which makes it possible to neglect a linear potential induced by the OFR field Banerjee ; Campbell . In this case, $\psi_{1,2}$ are the wave functions of the two species, the total energy of the binary BEC being:
$E=\int_{V}\left[\sum_{j}\left(\frac{\hbar^{2}|\nabla\psi_{j}|^{2}}{2m}+\frac{\mathcal{G}_{jj}}{2}|\psi_{j}|^{4}\right)+\mathcal{G}_{12}|\psi_{1}|^{2}|\psi_{2}|^{2}\right.$
$\phantom{=\;\;}\left.+\frac{8m^{3/2}}{15\pi^{2}\hbar^{3}}\left(\sum_{j}{\mathcal{G}_{jj}}|\psi_{j}|^{2}\right)^{5/2}\right]\ ,\ \ j=1,2.$ (3)
Here $\hbar$ and $m$ are the Planck’s constant and atomic mass, while $\mathcal{G}_{jj}=4\pi\hbar^{2}a_{jj}/{m}$ ($j=1,2$) and $\mathcal{G}_{12}=4\pi\hbar^{2}a_{12}/{m}$ are the strengths of the intra- and inter-component mean-field interactions, respectively. Here, $a_{11},a_{22}$ and $a_{12}$ are the corresponding $s$-wave scattering lengths of atomic collisions, and the last term of Eq. (3) accounts for the LHY correction Petrov2015 ; Fischer2006 ; Inguscio2018 .
For this system, the minimization of mean-field energy is attained at $n_{1}/n_{2}=\sqrt{a_{22}/a_{11}}$ (where $n_{1,2}=|\psi_{1,2}|^{2}$) Petrov2015 ; Inguscio2018 . Here, we consider the symmetric case with $a_{22}=a_{11}$, which leads to the condition of $n_{2}=n_{1}$, hence the two wave-function coalesce into a single one as $\psi_{1}=\psi_{2}\equiv\psi/\sqrt{2}$. We consider the setup with the inter-component scattering length engineered by means of OFR so as to set $a_{12}=a_{\mathrm{bg}}+a_{\mathrm{opt}}$, where $a_{\mathrm{bg}}<0$ is the background value of the scattering length in the absence of the laser illumination. It is adjusted by the additional magnetic Feshbach resonance as $a_{\mathrm{bg}}=-a_{11}$, to cancel the spatially uniform scattering length (the detailed consideration of the two-component asymmetry is presented in the Appendix A). Thus, energy (3) is simplified to
$E=\int_{V}\left(\frac{\hbar^{2}}{2m}|\nabla\psi|^{2}+\frac{\pi\hbar^{2}a_{\mathrm{opt}}}{m}|\psi|^{4}+\frac{2}{5}\Gamma|\psi|^{5}\right),$ (4)
where $\Gamma=4m^{3/2}\mathcal{G}_{11}^{5/2}/3\pi^{2}\hbar^{3}$. Accordingly, the dynamics is governed by the single Gross-Pitaevskii (GP) equation:
$i\hbar\frac{\partial\psi}{\partial t}=-\frac{\hbar^{2}\nabla^{2}}{2m}\psi-\frac{2\pi\hbar^{2}a^{(0)}}{m}e^{-\frac{\rho^{2}}{L^{2}}}|\psi|^{2}\psi+\Gamma|\psi|^{3}\psi.$ (5)
Stationary solutions of Eq. (5) with chemical potential $\mu$ and integer vorticity $S$ are looked for in the usual form,
$\psi\left(x,y,z,t\right)=\exp\left(-i\mu t+iS\theta\right)\phi\left(\rho,z\right),$ (6)
with angular coordinate $\theta$ and positive real function $\phi\left(\rho,z\right)$ satisfying the stationary equation
$\hbar\mu\phi=-\frac{\hbar^{2}}{2m}\nabla^{2}_{\rho z}\phi-\frac{2\pi\hbar^{2}a^{(0)}}{m}e^{-\frac{\rho^{2}}{L^{2}}}\phi^{3}+\Gamma\phi^{4}.$ (7)
[…]
III Characteristics of the Elongated states
To quantify the slenderness of 3D vortex QDs produced by Eq. (5), we define the aspect ratio, $\mathcal{A}=D_{\mathrm{v}}/D_{\mathrm{h}}$, where $D_{\mathrm{v}}$ and $D_{\mathrm{h}}$ are, respectively, vertical and horizontal lengths of the QDs. First, the numerical solution of the stationary version of Eq. (5) for fundamental QDs [with $S=0$ in Eq. (6)] demonstrates, in Figs. 1(a1-a3), the effect of the nonlinear trapping potential on the QD’s shape. In Fig. 1(a1), the aspect ratio is $\mathcal{A}=1$ for the isotropic fundamental QD in the free space ($L=\infty$), and $(a^{(0)},N)=(20a_{0},10^{6})$. As shown in Figs. 1(b1,b2), the aspect ratio increases (i.e., the fundamental QD elongates) with the decrease of radius $L$ of the holding beam in Eq. (2) and increase of $N$. The numerical solution of the time-dependent equation (5) (not shown here in detail) demonstrates that the entire family of the elongated QDs with $S=0$ is stable. Note, in particular, that the $\mu(N)$ dependence plotted in Fig. 1(b2) satisfies the well-known necessary stability condition, $d\mu/dN<0$ (the Vakhitov-Kolokolov criterion Vakhitov1973 ; Berge ).
Figure 1: (a1-a3) Typical examples of fundamental ($S=0$) QDs, composed of $(a^{(0)},N)=(20a_{0},10^{6})$ atoms and shaped by the OFR-induced spatial modulation (2) with $L=\infty$ (i.e., in the free space) (a1), $5\ \upmu\mathrm{m}$ (a2) and $2\ \upmu\mathrm{m}$ (a3). The vertical and horizontal lengths, $D_{v}$ and $D_{h}$, of the QDs are indicated in the panels. (b1) and (b2): The aspect ratio $\mathcal{A}$ and chemical potential $\mu$ of the QDs vs. $L$ and $N$, respectively.
In Fig. 1(b2), the chemical potential of the fundamental QDs initially drops steeply with the increase of $N$, and then, at $N\rightarrow\infty$, it saturates at an equilibrium value $\mu_{0}$. To qualitatively explain the saturation, one may take the 1D version of Eq. (7) for QDs strongly elongated in the direction of $z$, in which the transverse $\rho$-derivatives are dropped and $S=0$ is set:
$\hbar\mu\tilde{\phi}=-\frac{\hbar^{2}}{2m}\frac{d^{2}\tilde{\phi}}{dz^{2}}-\frac{2\pi\hbar^{2}a^{(0)}}{m}\tilde{\phi}^{3}+\Gamma\tilde{\phi}^{4}.$ (8)
This limit corresponds to reducing Eq. (7) for the strongly elongated QDs to the effectively 1D equation. It may be adopted as an approximation corresponding to the “quasi-inflexion point” quasi-inflexion . Further, a strongly elongated 1D solution to Eq. (8) seems as a long segment, $|z|<Z$, filled by the condensate with the nearly constant equilibrium density $\tilde{\phi}_{0}^{2}$, which is smoothly connected to the asymptotically zero solution, at $|z|>Z$, by kink solutions (alias fronts) around $z=\pm Z$ Birnbaum ; Petrov3 . To find the value of $\tilde{\phi}_{0}^{2}$ and the corresponding value $\tilde{\mu}_{0}$ of the chemical potential, one multiplies Eq. (8) by $d\tilde{\phi}/dz$ and performs the integration, reducing the second-order equation to the first-order one:
$\frac{1}{2}\hbar\mu\tilde{\phi}^{2}=-\frac{\hbar^{2}}{4m}\left(\frac{d\tilde{\phi}}{dz}\right)^{2}-\frac{\pi\hbar^{2}a^{(0)}}{2m}\tilde{\phi}^{4}+\frac{1}{5}\Gamma\tilde{\phi}^{5}.$ (9)
The kink solution to both equations (8) and (9), centered at $z=Z$, satisfy the boundary conditions $\phi^{2}(z-Z\rightarrow-\infty)=\tilde{\phi}_{0}^{2}$ and $\tilde{\phi}^{2}(z-Z\rightarrow+\infty)=0$. Because $z$-derivatives vanish at $z-Z\rightarrow\pm\infty$, in this limit Eqs. (8) and (9) amount to a system of algebraic equations for $\tilde{\phi}_{0}^{2}$ and $\tilde{\mu}_{0}$, which readily yields:
$\tilde{\phi}_{0}=\frac{5\hbar^{2}a^{(0)}}{3m\Gamma},~{}\tilde{\mu}_{0}=-\frac{25\hbar^{5}}{\Gamma^{2}}\left(\frac{\pi a^{(0)}}{3m}\right)^{3}.$ (10)
The comparison of the actual saturation value $\mu_{0}$ in Fig. 1(b2) with the one approximately predicted by Eq. (10) yields $\mu_{0}\approx 0.73\tilde{\mu}_{0}$, the discrepancy being explained by the omission of the transverse structure in the above consideration.
Figure 2: Characteristics of elongated vortex QDs with $S=1$. (a,b) Stable vortex QDs in the free space ($L=\infty$) (a) and shaped by the OFR-induced modulation with $L=2.5\ \upmu\mathrm{m}$ (b). (c-e) The aspect ratio $\mathcal{A}$ and chemical potential $\mu$ of the vortex QD, are plotted by the red and green curves (which correspond, respectively, to the left and right vertical axes) as functions $L$ (c), $N$ (d), $a^{(0)}$ (e), respectively. In these plots, solid and dashed lines represent stable and unstable solutions, respectively. (f) The stability chart for the vortex QDs with $S=1$ in the plane of ($a^{(0)},L$) with $N=10^{6}$.
Note that the elongated QDs with embedded integer vorticity $S$ also feature the saturation of $\mu$ at $N\rightarrow\infty$, see Fig. 4(d) below. As concerns the transverse structure of the strongly elongated QDs, it obviously corresponds to solutions of the 2D reduction of Eq. (7), produced by dropping the $z$ derivative in the equation.
Next, characteristics of elongated QDs with the embedded vorticity, $S=1$, are displayed in Fig. 2, illustrating the significant impact of varying parameters on their shape. The solution for the 3D vortex QD in the free space (it corresponds to $L=\infty$), which in plotted in Fig. 2(a), coincides with the one reported in Ref. Kartashov2018 . Figs. 2(b) show the shapes of the elongated vortex QDs when transverse confinement is applied. However, the vortex QDs are unstable when $N$ is insufficiently large, or the effective nonlinear potential is too weak, or the confinement is too tight (i.e., $a^{(0)}$ or $L$ is too small). In Fig. 2(c-e), the “slenderness” and stability of the vortex QDs with $S=1$ are systematically investigated with effects from $L$, $N$, and $a^{(0)}$. In Fig. 2(c), the decrease of $L$ significantly increases the aspect ratio, making the QD unstable when $L$ falls below a critical value. Fig. 2(d) shows that, with the increase of $N$, the QD becomes more elongated, with larger values of the aspect ratio. On the other hand, below a critical value of $N$, the vortex QDs are unstable. In Fig. 2(e), the variation of the modulation amplitude $a^{(0)}$ has minimal impact on the aspect ratio. The instability of the vortex QD occurs at $a^{(0)}$ falling below a critical value. The existence of the critical values implies that there is a stability boundary in the parametric space. Fig. 2(f) displays the stability area in the parameter plane of $(L,a^{(0)})$, which is produced by systematic simulations of the perturbed evolution of the vortex QD.
IV Dynamics
The independence of the modulation profile (2) on $z$ makes it possible to generate QDs freely moving along the $z$ axis, and thus simulate Eq. (5) for collisions between the QDs moving in opposite directions, initiated by the input, $\psi_{0}\left(\rho,z\right)=\psi(\rho,z-z_{d})\mathrm{e}^{-i\eta z}+\psi(\rho,z+z_{d})\mathrm{e}^{i\eta z}$, where $2z_{d}$ is the initial separation between the QDs, and ±$\pm\eta$ are the opposite kicks which set the droplets in motion. Figures 3(a1-a4) show that the collision between the QDs with identical vorticities $S=1$ is inelastic, resulting in the merger of the vortex QDs into a single oscillatory one (breather). The breathing dynamics is robust and quasi-harmonic, lasting, at least, for $>100\ \mathrm{ms}$, as is illustrated in Fig. 3(b) a sequence of $5$ cycles within the time interval of $50\thicksim 75\ \mathrm{ms}$. The breathing frequency is nearly independent of the collision velocity, suggesting that it is an eigenfrequency of a QD’s intrinsic mode. The collision leads to full destruction of the vortex QDs if the initial kick exceeds a certain critical value, which is $\eta_{\mathrm{cr}}=1.2\ \upmu\mathrm{m^{-1}}$. Collisions between zero-vorticity QDs produce similar results (not shown here in detail). On the other hand, for the QDs with opposite vorticities, $S=\pm 1$, the simulations demonstrate that the collision is quasi-elastic, leading to temporary fusion into a single droplet, which then splits back into original moving vortex QDs. The latter result is explained by the fact that the opposite vorticities attenuate the coherent interaction between the colliding objects. It is shown in Fig. 3(c) that the holding beam rotating around the $y$-axis in the $\left(xz\right)$-plane at angular velocity $\omega=0.163\ \mathrm{rad/ms}$, which means the replacement of the trapping profile (2) by
$a_{\mathrm{opt}}=-a^{(0)}\exp[-\frac{(x\cos(\omega t)-z\sin(\omega t))^{2}+y^{2}}{L^{2}}],$ (11)
is able to drive the adiabatic rotation of the elongated vortex QD. When the angular velocity $\omega$ exceeds the threshold value $0.977\ \mathrm{rad/ms}$, the enforced rotation causes the vortex droplet to split.
Figure 3: For all the droplets, $(L,a^{(0)},N)=(2.5\ \upmu\mathrm{m},20a_{0},5\times 10^{5})$. (a1-a4) Collision of identical elongated vortex QDs, with $S=1$, set in motion along the $z$ axis by kicks $\eta=1\ \upmu\mathrm{m^{-1}}$, the initial half-separation between the QDs being $z_{d}$ $=12.8\ \upmu\mathrm{m}$. Panel (b) illustrates quasi-harmonic oscillations of the breather produced by the merger of the colliding vortex QDs fitted to the sine function. (c) The rotation of an elongated vortex QD driven by the rotating trapping profile (11), with the resulting configurations displayed at different moments, which correspond to the rotation angles 00, $\pi/4$, and $\pi/2$, respectively. (d) The swing motion of the fundamental elongated QD in the $(xz)$-plane. (e1-e3) The gyroscopic motion (precession) of an elongated vortex QD, shown by projection of the center-of-mass trajectory onto the horizontal plane fixed at $z=2.5\ \upmu\mathrm{m}$. Panels (e1), (e2), and (e3) display the trajectories initiated by the torque with different strengths.
In addition, it is possible to excite swinging motion of an elongated QD in the stationary trapping profile (2) by applying an original torque to it. First, Fig. 3(d) shows that the torque, in the form of the multiplication of $\psi\left(x,y,z\right)$ by $\exp[i\alpha x\tanh\left({z/z_{0}}\right)]$ with strength $\alpha$, applied to a fundamental (zero-vorticity) elongated QD, sets it in the state of swinging motion in the $(xz)$-plane, with a period weakly depending on $\alpha$. In the case displayed in Fig. 3(d) the period is $0.792\ \mathrm{ms}$, for $(\alpha,z_{0})=(2\ \upmu\mathrm{m^{-1}},10\ \upmu\mathrm{m})$. Furthermore, the application of the torque to an elongated vortex QD initiates its gyroscopic motion (precession), cf. Ref. gyro . In the present case, this is done by suddenly tilting the vortex QD in the $(xz)$-plane. Then, it is either released or is additionally subjected to the action of the torque represented by factor $\exp[i\alpha y\tanh(z/z_{0})]$. Because the original vorticity is counter-clockwise with respect to the positive $z$-axis, this torque with $\alpha>0$ or $\alpha<0$ makes the vortex angular momentum smaller or larger, respectively. The precession initiated by this procedure is represented in Figs. 3(e1-e3) by the projection of trajectories of the QD’s center-of-mass onto the horizontal plane fixed at $z=2.5\ \upmu\mathrm{m}$, in the time interval $t=9\thicksim 90\ \mathrm{ms}$. In the case of zero torque ($\alpha=0$), Fig. 3(e1) exhibits an elliptic trajectory rotating around the center. For $\alpha>0$ and $\alpha<0$, Figs. 3(e2) and (e3) show that the trajectories become, respectively, more circular or more elliptical when the applied torque makes the resulting angular momentum smaller or larger, resulting the gyroscopic motion may be considered as a an example of macroscopic spin-orbit coupling.
Figure 4: (a-c) Typical examples of the vortex droplets and their phases diagrams (at $z=0$) with $(L,a^{(0)},N)=(8\ \upmu\mathrm{m},20a_{0},2\times 10^{6})$. (d) The chemical potential $\mu$ of the QDs with $S=0\sim 4$ vs. $N$, with solid and dashed lines representing stable and unstable solutions, respectively.
V Higher-order vortex states
The system modeled by Eq. (5) also supports stable QDs with higher values of the embedded vorticity (topological charge), $S=2$, $3$, and $4$, examples of which are displayed in Figs. 4(a-c). It is relevant to stress that this funding is an advantage of the nonlinear trapping potential, as the linear 3D harmonic-oscillator one can stabilize solely vortex solitons with $S=1$ Mihalache . Naturally, the increase of $S$ for fixed parameters $L$ and $a^{(0)}$ in the modulation profile (2), and fixed number of atoms $N$ makes the radius of the vortex’ inner hole larger, leading to a decrease in the aspect ratio $\mathcal{A}$. Families of the higher-order vortex QDs are characterized in Fig. 4(d) by respective dependences of the chemical potential $\mu$ on $N$, cf. similar dependences for $S=0$ and $1$ displayed in Figs. 1(b2) and 2(d), respectively. Naturally, Fig. 4(d) demonstrate that higher topological charges produce larger higher chemical potentials. This conclusion can be easily explained if one adopts the simple ansatz for the transverse shape of the vortex QD, $\phi\sim\rho^{S}\exp\left(-\rho^{2}/\rho_{0}^{2}\right)$. Indeed, substituting it in Eq. (7) one readily obtains the $S$-dependent shift of the chemical potential, $\delta\mu=\left(2\hbar/m\rho_{0}^{2}\right)S$. The respective splitting between the eigenvalues corresponding to vorticities $S$ and $S+1$, $\Delta\mu=2\hbar/m\rho_{0}^{2}$, very well matches the numerical results displayed in Fig. 4(d).
VI Conclusion
We have proposed to use the OFR, modulating the inter-component atomic $s$-wave scattering length in the binary BEC, to induce the effective 2D nonlinear trapping potential with the Gaussian profile. It supports elongated QDs, including ones with embedded integer vorticity $S$, which are stabilized by the GP-LHY equations. We have systematically explored the effects of the modulation amplitude and width, as well as the total number of atoms, on the shape and stability of the QDs. The fundamental QDs with $S=0$ are completely stable, while families of the vortex QDs with $S=1,2,3,$ and $4$ are partly stable. Collisions of identical elongated vortex QDs with $S=1$ lead to their merger into breathers with the same vorticity, $S=1$, while collisions between the elongated QDs with opposite vorticities, $S=\pm 1$, are quasi-elastic. The moving OFR modulation profile can rotate the elongated QDs in the abiabatic regime. The application of the torque to the trapped elongated vortex QD sets it in a state of stable gyroscopic motion (precession).
As an extension of the analysis, it may be interesting to consider more complex holding potentials, such as anisotropic ones, as well as those combining the nonlinear trap and vortex laser beams. Such modes with long inner voids can be used, in the context of the currently developing field matter-wave atomtronics, as conduits (waveguides) for linear atomic jets, similar to the concept of ”light guided by light”, which is wellknown in nonlinear optics Demircan ; Vasa ; Fang . Magnetic-gradient/radiofrequency-mediated scattering-length control, combined with droplet-creation schemes, could simplify future experiments Carli ; Sanz ; Lavoine .
Acknowledgements.
We appreciate valuable discussions with Jingxuan Sun. This work was supported by NNSFC (China) through Grants No. 12274077, No. 12475014, Guangdong Basic and Applied Basic Research Foundation 2023A1515110198, 2024A1515030131, 2025A1515011128, No. 2023A1515010770, the Research Fund of Guangdong-Hong Kong-Macao Joint Laboratory for Intelligent Micro-Nano Optoelectronic Technology through grant No.2020B1212030010. The work of B.A.M. was supported, in part, by the Israel Science Foundation through grant No. 1695/2022.
Appendix A Two-component asymmetry situation
According to the defination of the total energy (see Eq. (3) in the main text), the corresponding two-component GP-LHY equation is written as:
$i\hbar\frac{\partial\psi_{1}}{\partial t}=-\frac{\hbar^{2}}{2m}\nabla^{2}\psi_{1}+\left(\mathcal{G}_{11}|\psi_{1}|^{2}+\mathcal{G}_{12}|\psi_{2}|^{2}\right)\psi_{1}$
$+\gamma a_{11}\left(a_{11}|\psi_{1}|^{2}+a_{22}|\psi_{2}|^{2}\right)^{3/2}\psi_{1},$ (12)
$i\hbar\frac{\partial\psi_{2}}{\partial t}=-\frac{\hbar^{2}}{2m}\nabla^{2}\psi_{2}+\left(\mathcal{G}_{22}|\psi_{2}|^{2}+\mathcal{G}_{12}|\psi_{1}|^{2}\right)\psi_{2}$
$+\gamma a_{22}\left(a_{11}|\psi_{1}|^{2}+a_{22}|\psi_{2}|^{2}\right)^{3/2}\psi_{2},$ (13)
where $\gamma=128\sqrt{\pi}\hbar^{2}/3m$, and we consider the setup with the inter-component scattering length engineered by means of OFR as $a_{12}=a_{\mathrm{bg}}+a_{\mathrm{opt}}=a_{\mathrm{bg}}-a^{\left(0\right)}\exp\left(-\rho^{2}/L^{2}\right)$ (where $\rho\equiv\sqrt{x^{2}+y^{2}}$). Here, $a_{\mathrm{bg}}<0$ is the background value of the scattering length in the absence of light, which is adjusted by the additional magnetic Feshbach resonance and characterizes an attractive interaction between the two species of $\psi_{1}$ and $\psi_{2}$. Further, $L$ is the beam waist, and $-a^{\left(0\right)}<0$ is the amplitude of the OFR-induced modulation of the local scattering length, the total numbers of atoms in the two components are
[…]
Regarding the asymmetric situation discussed here, under the requirement of the minimization of the mean-field energy, the ratio of the equilibrium densities is locked as $n_{1}/n_{2}=\sqrt{\mathcal{G}_{22}/\mathcal{G}_{11}}$ Petrov2015 , hence the ratio of the particle numbers is also $N_{1}/N_{2}=\sqrt{a_{22}/a_{11}}$ Inguscio2018 , which results in $a_{22}=\left(N_{1}/N_{2}\right)^{2}a_{11}$. Here we keep fixing $a_{\mathrm{bg}}=-a_{11}$ and simulate the elongated vortex with the imbalanced density by adjusting the total number of atoms $N_{1,2}$ in each of the two components.
Figure 5: Typical examples of the asymmetric population imbalance for the two-components system, with $N_{1}=2.4\times 10^{5}$, $N_{2}=2.6\times 10^{5}$, $L=2.5\mathrm{\mu m}$, $a^{(0)}=20a_{0}$, and $a_{11}=50a_{0}$. (a) and (b) represent the component $\psi_{1}$ and $\psi_{2}$, respectively.
First, we consider the situation with slight imbalances. As shown in Fig. (5), we set $N_{1}=2.4\times 10^{5},N_{2}=2.6\times 10^{5}$, we have constructed the corresponding imbalanced slender vortex solution and demonstrated its stability through direct evolution.
Then, we consider the situation with slight imbalances. We set $N_{1}=2\times 10^{5},N_{2}=3\times 10^{5}$, the corresponding imbalanced slender vortex solution is unstable, the direct evolution as shown in Fig. (6)
Figure 6: Typical examples and evolution of the unstable solution, with $N_{1}=2\times 10^{5}$, $N_{2}=3\times 10^{5}$, $L=2.5\mathrm{\mu m}$, $a^{(0)}=20a_{0}$, and $a_{11}=50a_{0}$. (a1-a3) and (b1-b3) represent the component $\psi_{1}$ and $\psi_{2}$, respectively.
The above results show that the size profiles of the two components are the same, but the densities are different. Overall, we have found that the stable vortices persist if the population imbalance between the two components is not too large. On the other hand, if the intercomponent population imbalance attains a critical value, the vortex QD becomes unstable and decays to the ground state within a few microseconds.